# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v1

@@ 1 Introduction
where $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})$ and independently, $\varepsilon_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2})$ for all $t$. We denote the differential parameter by $\bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0}\in\mathbb{R}^{p}$, the maximum variability of the response by $\Psi=\max(|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2},|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2},\sigma)$, and the minimal segment length by $\Delta=\min(\theta,n-\theta)$ As a convention, we write $\theta=n$, when $\bm{\beta}_{0}=\bm{\beta}_{1}\in\mathbb{R}^{p}$, corresponding to the case of no change. By focusing on the AMOC situation under (1), we separate the two “orthogonal” aspects of the problem: The design of an adaptively optimal procedure for change point detection and estimation, which is the main focus of this paper, and the development of algorithms for multiple change point detection. While our theoretical analysis is framed under independence and Gaussianity, we also discuss extensions to accommodate temporal dependence or non-Gaussianity.
@@
    Phase transition and a new notion of sparsity. We derive minimax lower bounds for both detection and estimation of the change point, as summarised in Table 1. The (nearly) matching upper and lower bounds presented therein reveal a phase transition governed by $\mathfrak{s}:=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$, which we term the inherent sparsity, rather than the standard notion of sparsity $\mathfrak{s}_{\delta}:=|\bm{\delta}|_{0}$, where $|\cdot|_{0}$ denotes the $\ell_{0}$-norm. Notably, the results also indicate that in the dense regime, where $\mathfrak{s}\geq\sqrt{p}$, consistent estimation is possible only when $p\lesssim n^{2}$. See Section 2 for details.
@@
    $\mathfrak{s}<\sqrt{p}$ ${\mathfrak{s}}\log(\tfrac{ep}{\mathfrak{s}^{2}})$ $\sigma_{*}^{-1}\mathfrak{s}\log(p)$ $\frac{{\Psi^{2}}}{|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\mathfrak{s}\log(\tfrac{ep}{\mathfrak{s}^{2}})$ $\frac{\Psi^{2}}{\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\mathfrak{s}\log(p)$ McScan (Section 3.1)
    $\mathfrak{s}\geq\sqrt{p}$ $\sqrt{p}$ $\sigma_{*}^{-1}\sqrt{p\log\log(n)}$ $\frac{{\Psi^{2}}}{|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\sqrt{p}$ $\frac{\Psi^{2}}{\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\sqrt{p\log\log(n)}$ QcScan (Section 3.2)
@@
    The first method for consistent change point detection in the dense regime. In a similar spirit to McScan (Cho et al.,, 2025), which is shown to be minimax optimal in the sparse regime (Section 3.1), we introduce the quadratic covariance scanning (QcScan) that evaluates the $\ell_{2}$-aggregation of the contrasts in local sample covariances between regressors and responses. QcScan is the first method that achieves minimax optimality, up to a factor of $\sqrt{\log\log(n)}$, in detection and estimation of the change point in the dense regime. As a by-product, we also obtain an estimator for $|\bm{\Sigma}\bm{\delta}|_{2}^{2}$, which serves as a surrogate of the strength of the signal $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ for the change point problem in (1). See Section 3.2 for details.
@@
    Adaptation to inherent sparsity. We introduce OcScan which, by combining McScan and QcScan, adapts to the unknown level of sparsity and thus achieves optimal statistical performance uniformly across all $\mathfrak{s}$, providing a complete treatment of the change point detection problem in high-dimensional linear models. See Section 4.1 for details.
@@ 2.1 Isotropic case
Let $\sigma>0$ and $41\leq\Delta\leq n/$ 4, and suppose that $\bm{\Sigma}=\mathbf{I}$.
@@ 2.2 General covariance
Figure 2 numerically confirms these observations when $\mathsf{Cov}(\mathbf{x}_{t})=[0.6^{|i-j|}]_{i,j}$. Specifically, the best performance of various change point estimation methods exhibits a steeper dependence on the intrinsic sparsity (middle panel) than on the standard sparsity (left panel), aligning well with the theoretical bounds shown in Table 1 (right panel). Besides, we notice that McScan, a method attaining minimax rate-optimality in the sparse regime, has its estimation error exceed that of QcScan, a method tailored to achieve optimal performance in the dense regime, close to where the inherent sparsity $\mathfrak{s}$ meets $\sqrt{p}$ (middle panel). On the other hand, this phenomenon occurs far before the sparsity of $\bm{\delta}$ reaches $\mathfrak{s}_{\delta}=\sqrt{p}$, which is accounted for by that $|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=p$ in all the scenarios in consideration in the left panel.
Figure 2: Estimation performance of MOSEG (Cho and Owens,, 2024), CHARCOAL (Gao and Wang,, 2022), and the proposed methods: McScan (Section 3.1; a variant of Cho et al.,, 2025), QcScan (Section 3.2), OcScan (Section 4.1) and OcScan.R (Section 4.2), when $\bm{\Sigma}=[0.6^{|i-j|}]_{i,j}$. Taken from (M2) in Section 5, we set $n=300$, $p=200$, $\bm{\beta}_{0}+\bm{\beta}_{1}=\mathbf{0}$ and $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}=2$. In the left, the $x$-axis denotes the standard sparsity $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$ while in the middle, it is the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. In each scenario, results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The right panel plots the quantity $|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}\wedge\sqrt{p}$ as either $\mathfrak{s}_{\delta}$ or $\mathfrak{s}$ varies. This quantity is the influencing factor, save logarithmic factors, appearing in the minimax rate of estimation (see Table 1), and represents a proxy of the difficulty of the estimation problem. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$ or $\mathfrak{s}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@ 3.1 Sparse regime
Let $\sigma_{*}$ denote the smallest non-zero eigenvalue of $\bm{\Sigma}$. By Hölder’s inequality, we have
@@
Thus, as long as $\sigma_{*}$ is bounded away from zero, McScan is minimax rate optimal up to a logarithmic factor in the sparse regime $\mathfrak{s}<\sqrt{p}$ (cf. Table 1), without requiring any prior knowledge of $\bm{\Sigma}$, or ruling out strong collinearity among the variables that results in rank deficient $\bm{\Sigma}$. When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ as in (2) allows for re-formulating the condition in (5), to
@@
The original McScan algorithm of Cho et al., (2025) combines the detector statistic in (3) with the narrowest-over-threshold principle (Baranowski et al.,, 2019) and the seeded intervals (Kovács et al.,, 2023), evaluating the detector statistics over strategically selected intervals and selecting the significant local maximisers for the detection of multiple change points. Theorem 3 is a variant of the more general Theorem 2 in Cho et al., (2025) which, permitting serial dependence and sub-Weibull tail behaviour in the data, shows the consistency of McScan in multiple change point detection, all without placing any condition on the exact sparsity of regression or differential parameters, or that on $p$ beyond what is implied by the detection lower bound comparable to (5). There, it is further demonstrated that McScan allows for a strictly wider class of change point models in comparison with the state-of-the-art, see Section S1 in the supplementary material of Cho et al., (2025).
@@ 3.2 Dense regime
Thus, QcScan is minimax near-optimal up to a multiplicative factor of $\sqrt{\log\log(n)}$, in the dense regime where $\mathfrak{s}\geq\sqrt{p}$ (cf. Table 1), provided that $\sigma_{*}$ is bounded away from zero. When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ allows us to re-write (8) as
$|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta\geq c_{0}\Psi^{2}\sqrt{p\log\log(n)},$
which removes the dependence on $\sigma_{*}$.
@@ 3.3.1 When the differential parameter is sparse
In the inherently sparse regime with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny M}}$, the consistency of $\widehat{\bm{\delta}}$ is analogous to that of the Lasso estimator of the regression parameter (see e.g. Bickel et al.,, 2009), with $\Delta=\min(\theta,n-\theta)$ taking the role of sample size. Cho et al., (2025) additionally propose a bias-corrected estimator of $\bm{\delta}$ and establish a non-asymptotic Gaussian approximation result, which enables simultaneous inference about all $p$ coordinates of $\bm{\delta}$. Later in Section 4.2, we devise a change point estimator which, based on scanning the projected covariance as $\widehat{\bm{\delta}}^{\top}\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)$, attains a refined rate of estimation that is inversely proportional to $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$.
@@ 4.1 Adaptively optimal change point detection
with $\text{with \ }C_{\text{\tiny M}/\text{\tiny Q}}=\left(\frac{T_{\widehat{\theta}_{\text{\tiny Q}}}}{\sqrt{p\log\log(n)}}\right)^{-1}\frac{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}^{2}}{\log(p\log(n))},$
@@
The change point problem in linear models can be viewed as a special case of the change point problem in covariances, which allows us to place the performance of our proposed methods in a context. This is particularly fruitful in elucidating the role of terms such as $\|\bm{\Sigma}\|$ and $\Psi$ that are often treated as constants in the literature on change point detection in linear models. Compared to the state-of-the-art methods (Wang et al., 2021a, ; Dette et al.,, 2022; Moen,, 2024) that target different types of changes in covariances, our OcScan consistently achieves equal or superior performance across a variety of structural settings. For instance, OcScan attains a sharper detection boundary than the projection-based method of Wang et al., 2021a , with an improvement factor of $p\min(\|\bm{\Sigma}\|\sigma_{*}^{-1}\mathfrak{s},|\bm{\Sigma}\bm{\delta}|_{0},\sqrt{p})^{-1}$. This gain arises from effectively leveraging the specific structure inherent in the linear model. Detailed discussion can be found in Appendix A.
@@ 4.2 Refined change point estimation
The rate reported in Proposition 9 is comparable to the rate obtained in Theorem 3 of Xu et al., (2024). Their change point estimator attains the rate $O_{P}(|\bm{\delta}|_{2}^{-2})$, which allows for the derivation of its limiting distribution, provided that $|\bm{\delta}|_{2}^{2}\Delta\gtrsim\mathfrak{s}_{\delta}^{2}\log^{3}(p\vee n)$, $\bm{\Sigma}$ have bounded eigenvalues (i.e. $\underline{\sigma}^{-1}\|\bm{\Sigma}\|$ is fixed) and $\Psi\lesssim 1$. Under comparable conditions, from that $|\bm{\delta}|_{2}^{-2}\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}\geq\underline{\sigma}$, the requirement in (12) is weaker by the factor of $\log^{2}(p\vee n)$, and clearly specifies the effects of $\|\bm{\Sigma}\|$, $\underline{\sigma}$ and $\Psi$. We also remark that Lemma 2 shows the inverse of $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ to be a better measure of the inherent difficulty of the detection problem in comparison with $|\bm{\delta}|_{2}^{-2}$.

### v2

@@ 1 Introduction
where $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})$ with non-negative definite $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ and independently, $\varepsilon_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2})$ for all $t$. We denote the differential parameter by $\bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0}\in\mathbb{R}^{p}$, the maximum variability of the response by $\Psi=\max(|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2},|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2},\sigma)$, and the minimal segment length by $\Delta=\min(\theta,n-\theta)$ As a convention, we write $\theta=n$, when $\bm{\beta}_{0}=\bm{\beta}_{1}\in\mathbb{R}^{p}$, corresponding to the case of no change. By focusing on the AMOC situation under (1), we separate the two “orthogonal” aspects of the problem: The design of an adaptively optimal procedure for change point detection and estimation, which is the main focus of this paper, and the development of algorithms for multiple change point detection. While our theoretical analysis is framed under independence and Gaussianity, we also discuss extensions to accommodate temporal dependence or non-Gaussianity.
@@
    Phase transition and a new notion of sparsity. We derive minimax lower bounds for both detection and estimation of the change point, as summarised in Table 1. The (nearly) matching upper and lower bounds presented therein reveal a phase transition governed by $\mathfrak{s}:=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$, which we term the inherent sparsity, rather than the standard notion of sparsity $\mathfrak{s}_{\delta}:=|\bm{\delta}|_{0}$, where $|\cdot|_{0}$ denotes the $\ell_{0}$-norm. Notably, the results also indicate that in the dense regime, where $\mathfrak{s}\geq\sqrt{p}$, consistent estimation is possible only when $p\lesssim n^{2}$. See Section 2.
@@
    $\mathfrak{s}<\sqrt{p}$
    ${\mathfrak{s}}\log(\tfrac{ep}{\mathfrak{s}^{2}})$
    $\sigma_{*}^{-1}\mathfrak{s}\log(p)$
    $\frac{{\Psi^{2}}}{|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\mathfrak{s}\log(\tfrac{ep}{\mathfrak{s}^{2}})$ $\frac{\Psi^{2}}{\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\mathfrak{s}\log(p)$
    McScan
    (Section 3.1)
    $\mathfrak{s}\geq\sqrt{p}$
    $\sqrt{p}$
    $\sigma_{*}^{-1}\sqrt{p\log\log(n)}$
    $\frac{{\Psi^{2}}}{|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\sqrt{p}$ $\frac{\Psi^{2}}{\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\sqrt{p\log\log(n)}$
    QcScan
    (Section 3.2)
@@
    The first method for consistent change point detection in the dense regime. We first show that McScan (Cho et al.,, 2025), which scans the $\ell_{\infty}$-aggregation of the contrasts in local sample covariance between regressors and responses, is minimax optimal in the sparse regime up to the factor of $\sigma_{*}^{-1}$ (Section 3.1), where $\sigma_{*}$ denotes the smallest non-zero eigenvalue of $\bm{\Sigma}$, i.e. McScan achieves optimality in the sparse regime even when $\bm{\Sigma}$ is rank deficient. For the dense regime, we introduce the quadratic covariance scanning (QcScan) that evaluates the $\ell_{2}$-aggregation of the local sample covariance contrasts. QcScan is the first method that achieves minimax optimality, up to a factor of $\sigma_{*}^{-1}\sqrt{\log\log(n)}$, in detection and estimation of the change point in the dense regime even when $\bm{\Sigma}$ is rank deficient, which is in contrast to existing change point methods that require $\mathsf{Cov}(\mathbf{x}_{t})$ to have bounded spectrum. As a by-product, we also obtain an estimator for $|\bm{\Sigma}\bm{\delta}|_{2}^{2}$, which serves as a surrogate of the strength of the signal $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ for the change point problem in (1). See Section 3.2.
@@
    Adaptation to inherent sparsity. We introduce OcScan which, by combining McScan and QcScan, adapts to the unknown level of sparsity and thus achieves optimal statistical performance uniformly across all $\mathfrak{s}$, providing a complete treatment of the change point detection problem in high-dimensional linear models. See Section 4.1.
@@ 2.1 Isotropic case
Let $\sigma>0$ and $1\leq\Delta\leq n/4$, and suppose that $\bm{\Sigma}=\mathbf{I}$.
@@ 2.2 General covariance
Figure 2 numerically confirms these observations when $\mathsf{Cov}(\mathbf{x}_{t})=[0.6^{|i-j|}]_{i,j}$. Specifically, the best performance of various change point estimation methods exhibits a steeper dependence on the intrinsic sparsity (middle panel) than on the standard sparsity (left panel), which is in alignment with the theoretical bounds shown in Table 1 (see also the right panel). Besides, we notice that McScan, a method attaining minimax rate-optimality in the sparse regime, has its estimation error exceed that of QcScan, a method tailored to achieve optimal performance in the dense regime, close to where the inherent sparsity $\mathfrak{s}$ meets $\sqrt{p}$ (middle panel). On the other hand, this phenomenon occurs far before the sparsity of $\bm{\delta}$ reaches $\mathfrak{s}_{\delta}=\sqrt{p}$, which is accounted for by that $|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=p$ in all the scenarios considered in the left panel.
Figure 2: Estimation performance of MOSEG (Cho and Owens,, 2024), CHARCOAL (Gao and Wang,, 2022), and the proposed methods: McScan (Section 3.1; a variant of Cho et al.,, 2025), QcScan (Section 3.2), OcScan (Section 4.1) and OcScan.R (Section 4.2), when $\bm{\Sigma}=[0.6^{|i-j|}]_{i,j}$. Taken from (M2) in Section 5, we set $n=300$, $p=200$, $\bm{\beta}_{0}+\bm{\beta}_{1}=\mathbf{0}$ and $|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}=2$. In the left, the $x$-axis denotes the standard sparsity $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$ while in the middle, it is the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. In each scenario, results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The right panel plots the quantity $\min(|\bm{\Sigma}^{1/2}\bm{\delta}|_{0},\sqrt{p})$ as either $\mathfrak{s}_{\delta}$ or $\mathfrak{s}$ varies. This quantity is the influencing factor, save logarithmic factors, appearing in the minimax rate of estimation (see Table 1), and represents a proxy of the difficulty of the estimation problem. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$ or $\mathfrak{s}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@ 3.1 Sparse regime
With $\sigma_{*}$ denoting the smallest non-zero eigenvalue of $\bm{\Sigma}$, we have by Hölder’s inequality,
@@
Thus, as long as $\sigma_{*}$ is bounded away from zero, McScan is minimax rate optimal up to a logarithmic factor in the sparse regime $\mathfrak{s}<\sqrt{p}$ (cf. Table 1), without requiring any prior knowledge of $\bm{\Sigma}$, or ruling out strong collinearity among the variables that results in rank deficient $\bm{\Sigma}$ – a situation that often arises with increasing number of variables. This is in stark contrast to the most existing methods that require $\mathsf{Cov}(\mathbf{x}_{t})$ to have bounded spectrum or characterise their properties in terms of the smallest eigenvalue of $\bm{\Sigma}$; see Section S1.1 of Cho et al., (2025) and also Figure 5 in Section 5 for numerical validation. When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ as in (2) allows for re-formulating the condition in (5), to
@@
The original McScan algorithm of Cho et al., (2025) combines the detector statistic in (3) with the narrowest-over-threshold principle (Baranowski et al.,, 2019) and the seeded intervals (Kovács et al.,, 2023), evaluating the detector statistics over strategically selected intervals and selecting the significant local maximisers for the detection of multiple change points. Theorem 3 is a variant of the more general Theorem 2 in Cho et al., (2025) which, permitting serial dependence and sub-Weibull tail behaviour in the data, shows the consistency of McScan in multiple change point detection, all without placing any condition on the exact sparsity of regression or differential parameters, or that on $p$ beyond what is implied by the detection lower bound comparable to (5). There, it is further demonstrated that McScan allows for a strictly wider class of change point models in comparison with the state-of-the-art.
@@ 3.2 Dense regime
Thus, QcScan is minimax near-optimal up to a multiplicative factor of $\sqrt{\log\log(n)}$, in the dense regime where $\mathfrak{s}\geq\sqrt{p}$ (cf. Table 1), provided that $\sigma_{*}$ is bounded away from zero (see also the discussions below Theorem 3). When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ allows us to re-write (8) without the dependence on $\sigma_{*}$, as
$|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta\geq c_{0}\Psi^{2}\sqrt{p\log\log(n)}.$
@@ 3.3.1 When the differential parameter is sparse
For the consistency of $\widehat{\bm{\delta}}$, we make the stronger assumption requiring $\bar{\sigma}>0$, compared to that required for the consistency in change point detection involving $\sigma_{*}$, the smallest non-zero eigenvalue of $\bm{\Sigma}$. In the inherently sparse regime with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny M}}$, the consistency of $\widehat{\bm{\delta}}$ is analogous to that of the Lasso estimator of the regression parameter (see e.g. Bickel et al.,, 2009), with $\Delta=\min(\theta,n-\theta)$ taking the role of sample size. Cho et al., (2025) additionally propose a bias-corrected estimator of $\bm{\delta}$ and establish a non-asymptotic Gaussian approximation result, which enables simultaneous inference about all $p$ coordinates of $\bm{\delta}$. Later in Section 4.2, we devise a change point estimator which, based on scanning the projected covariance as $\widehat{\bm{\delta}}^{\top}\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)$, attains a refined rate of estimation that is inversely proportional to $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$.
@@ 4.1 Adaptively optimal change point detection
$\text{with \ }C_{\text{\tiny M}/\text{\tiny Q}}=\left(\frac{T_{\widehat{\theta}_{\text{\tiny Q}}}}{\sqrt{p\log\log(n)}}\right)^{-1}\frac{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}^{2}}{\log(p\log(n))},$
@@
The change point problem in linear models can be viewed as a special case of the change point problem in covariances, which allows us to place the performance of our proposed methods in a context. This is particularly fruitful in elucidating the role of terms such as $\|\bm{\Sigma}\|$ and $|\bm{\Sigma}^{1/2}\bm{\beta}_{j}|_{2}$ (through $\Psi$) that are often treated as constants in the literature on change point detection in linear models. Compared to the state-of-the-art methods (Wang et al., 2021a, ; Dette et al.,, 2022; Moen,, 2024) that target different types of changes in covariances, our OcScan consistently achieves equal or superior performance across a variety of structural settings. For instance, OcScan attains a sharper detection boundary than the projection-based method of Wang et al., 2021a , with an improvement factor of $p\min(\|\bm{\Sigma}\|\sigma_{*}^{-1}\mathfrak{s},|\bm{\Sigma}\bm{\delta}|_{0},\sqrt{p})^{-1}$. This gain arises from effectively leveraging the specific structure inherent in the linear model. Detailed discussion can be found in Appendix A.
@@ 4.2 Refined change point estimation
The rate reported in Proposition 9 is comparable to the rate obtained in Theorem 3 of Xu et al., (2024). Their change point estimator attains the rate $O_{P}(|\bm{\delta}|_{2}^{-2})$, which allows for the derivation of its limiting distribution, provided that $|\bm{\delta}|_{2}^{2}\Delta\gtrsim\mathfrak{s}_{\delta}^{2}\log^{3}(p\vee n)$, $\bm{\Sigma}$ has bounded eigenvalues (i.e. $\underline{\sigma}^{-1}\|\bm{\Sigma}\|$ is fixed) and $\Psi\lesssim 1$. Under comparable conditions, from that $|\bm{\delta}|_{2}^{-2}\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}\geq\underline{\sigma}$, the requirement in (12) is weaker by the factor of $\log^{2}(p\vee n)$, and clearly specifies the effects of $\|\bm{\Sigma}\|$, $\underline{\sigma}$ and $\Psi$. We also remark that Lemma 2 shows the inverse of $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ to be a better measure of the inherent difficulty of the detection problem in comparison with $|\bm{\delta}|_{2}^{-2}$.